OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. y = -x + 6 y = x - 6 y = -x - 2 y = x + 2

OpenStudy (anonymous):

so to find a line that is perpendicular, we need to take the slope from this equation, y = -x + 6., turn it into a fraction, flip the fraction upside down and then times it by -1 can you try doing this for me?

OpenStudy (anonymous):

Would the fraction be 1/6?

OpenStudy (anonymous):

close! so the slope is found by taking only the coefficient of the 'x' term

OpenStudy (anonymous):

in our case it is -1 so with the slope of -1 turn it into a fraction, flip the fraction upside down and then times it by -1

OpenStudy (anonymous):

ok, so I would turn -1 into a fraction?

OpenStudy (anonymous):

yep ^_^

OpenStudy (anonymous):

Hmm so would that be -1 / 6?

OpenStudy (anonymous):

Ohh wait, would it look something like y + 2 = 1/2 (x + 4)?

OpenStudy (anonymous):